# PainDiffusion: Can robot express pain?

### v1

@@ Abstract
Pain is a more intuitive and user-friendly way of communicating problems, making it especially useful in rehabilitation nurse training robots. While most previous methods have focused on classifying or recognizing pain expressions, these approaches often result in unnatural, jiggling robot faces. We introduce PainDiffusion, a model that generates facial expressions in response to pain stimuli, with controllable pain expressiveness and emotion status. PainDiffusion leverages diffusion forcing to roll out predictions over arbitrary lengths using a conditioned temporal U-Net. It operates as a latent diffusion model within EMOCA’s facial expression latent space, ensuring a compact data representation and quick rendering time. For training data, we process the BioVid Heatpain Database, extracting expression codes and subject identity configurations. We also propose a novel set of metrics to evaluate pain expressions, focusing on expressiveness, diversity, and the appropriateness of model-generated outputs. Finally, we demonstrate that PainDiffusion outperforms the autoregressive method, both qualitatively and quantitatively. Code, videos, and further analysis are available at: https://damtien444.github.io/paindf/.
@@ I INTRODUCTION
Pain is defined as ”an unpleasant sensory and emotional experience associated with actual or potential tissue damage.” The perception and expression of pain vary among individuals due to their unique life experiences [1]. In this work, we propose utilizing pain expressions in robots as a means of communicating system problems. Traditional error codes, such as a red number on a screen, can be obscure, easily misunderstood, and often require specialized knowledge to interpret. In contrast, pain expressions provide a more intuitive and universal method for indicating issues, enabling users of all backgrounds to understand and respond appropriately. Much like how animals, infants, and unconscious beings express distress, incorporating pain-like signals into robots creates a more user-friendly interface. Pain facial expressions can facilitate the immediate recognition of problems, which helps protect the system from further damage due to improper handling or adverse working conditions. By leveraging the inherent human understanding of pain, our approach aims to improve the interaction between users and robotic systems, ensuring timely and effective responses to potential issues.
Beyond this application, rehabilitation training robots must also express pain as feedback to nursing learners. During rehabilitation exercises, nurses must be able to read patients’ emotions to improve the quality of care. Previous works confirm that recognizing patients’ pain from their facial behavior enhances nursing skills and overall care quality [2, 3, 4]. Therefore, incorporating pain facial expressions in robotic patients can be valuable for nurse and doctor training, ultimately improving healthcare standards, especially for the subtle facial expressions that might help treatment more ease for patients.
In this work, our contribution is threefold. First, we propose PainDiffusion, a model designed to generate pain facial expressions that can handle arbitrary-length predictions, making it suitable for robotic applications. Second, we incorporate intrinsic characteristics, such as expressiveness and emotion, allowing for a more controllable generation tailored to diverse use cases. Third, we introduce a new set of baselines and metrics to effectively evaluate the quality and accuracy of pain expressions generated by our model.
@@
Figure 1: Overview. PainDiffusion inputs pain stimuli signals, expressiveness configuration, emotion status, and past frames to generate the next appropriate prediction. By utilizing EMOCA latent as the facial expression representative, we effectively limit the data space, ensuring a smaller training time and inference time.
@@ II RELATED WORKS
In rehabilitation nurse-training robots, earlier work primarily focused on recognizing pain situations and selecting expressions from a predefined set [4, 5]. However, this method often results in expressions that are unnatural and less effective. More recently, research has shifted towards using generative models for facial expressions [6, 7]. While non-likelihood generative models like Generative Adversarial Networks (GANs) have produced high-quality results across various domains, they face significant challenges due to the complexity and instability of training and optimization. Meanwhile, diffusion models and score-based models have emerged as powerful tools for generating images and videos. These models have now achieved state-of-the-art performance, particularly in generating realistic images and videos [8, 9, 10]. In human action generation, recent work is increasingly focusing on diffusion models to achieve state-of-the-art results [11, 12, 13].
Among pain-related research, most efforts have focused on recognition and classification using either traditional machine learning or deep learning techniques [14, 15, 16, 17]. PainDiffusion is the first method to explore pain expression generation using diffusion models, supporting arbitrary-length predictions.
@@ III PAIN DIFFUSION
Our main goal is to model the relationship between facial expressions, pain-causing signals, and the intrinsic features of an individual, similar to the task of generating facial reactions based on a speaker’s speech or audio [6, 18]. To achieve this, we define the high-level task as follows: Given a continuous sequence of pain-causing signals and the configuration of the individual, we autoregressively predict the appropriate facial expression.
@@ III-C Temporal Latent Unet
Latent diffusion models (LDMs) have been effectively utilized for generating images, videos, and human behaviors [23, 10, 13, 8, 24]. In our work, we follow the LDM approach by restricting data representation to the EMOCA latent space, resulting in a much smaller data space compared to conventional diffusion models for image or video generation. Prior works in video generation diffusion models [25, 23] have introduced temporal layers to the standard spatial U-Net [26] to better capture temporal information. Instead of employing 2D convolutions as in [10], we hypothesize that our facial latent space does not retain substantial structural information of $I$ after multiple layers of encoding in EMOCA [19]. Therefore, using 1D convolutions is sufficiently effective.
The network is designed with convolutional ResNet blocks [27], each followed by a spatial attention block and then a temporal attention block. The noise $\lambda_{t}$ is integrated into the ResNet block using scale and shift operators, while conditioning information $c=(C,\mathcal{P},\mathcal{E})$ is encoded with a lightweight MLP encoder and incorporated into the spatial attention block via cross-attention. To embed temporal information, we scale and shift the features using temporal embeddings. Both the temporal and noise-time embeddings are encoded using sinusoidal embeddings [28] for positional information. The up and down blocks have the architecture as illustrated in Fig. 2.
@@ III-D Elucidated Diffusion
where $\boldsymbol{z}$ is the facial latent vector, $\sigma$ is the standard deviation of the Gaussian noise level, $C$ represents conditions, and $F_{\theta}$ is the temporal latent U-Net. By adjusting the terms $c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$, different diffusion strategies can be achieved with minimal changes to the model. We adopt the $c$ terms from the Elucidated Diffusion Model (EDM) [30] due to the flexibility of the network architecture. As a result, we do not consider reparameterization approaches, such as predicting velocity [31], since EDM [30] is a hybrid of both velocity, start, and noise prediction [32].
For clarity, the training objective is simplified as:
@@ III-E Diffusion Forcing
The goal of this model is to be applied to a robot and generate arbitrary-length predictions. We adopt diffusion forcing [29] to extend beyond the short training horizon. Diffusion forcing assigns different noise levels to each temporal timestep (Fig. 3), placing more uncertainty on future frames while reducing uncertainty for past frames. This allows the model to generate arbitrary-length predictions without divergence, as seen in [23]. Additionally, diffusion forcing eliminates the need for special guidance or extra forward passes to ensure temporal coherence, as required in [25]. It also introduces a hyperparameter that controls uncertainty—higher values result in greater uncertainty, which in turn requires more denoising steps.
Figure 3: Diffusion Forcing and Full Sequence Diffusion. Diffusion Forcing (a) applies different noise levels to each frame, whereas Full Sequence Diffusion (b) uses the same noise level for all frames.
@@
Figure 4: Qualitative comparison with Ground truth and Autoregressive baseline. Please visit the project website for better video qualitative results and more samples.
@@ IV EXPERIMENTS
Our experiments are designed with two primary goals. First, we aim to demonstrate that our model outperforms the baseline and common approaches in facial expression generation. Second, we seek to show that our model can produce arbitrary-length predictions without divergence. To achieve these objectives, we propose a new set of metrics specifically for evaluating pain facial expression and long-term prediction errors.
@@ IV-A Dataset
To create the pain facial expression dataset, we use the BioVid Heatpain Database part C [35, 36, 37], which captures heat pain responses from 87 subjects with 4 levels of pain intensity and 30 minutes of data, separated by pauses. The original dataset includes 3 modalities: frontal face video (25 FPS), biomedical signals, and heat stimuli signals. We split the dataset into two subsets: 61 subjects for training and 26 subjects for validation. The validation subsets maintain an equal ratio of male and female participants, and include 5 low-expression and 21 normal-expression subjects. We filter the original validation videos to contain sequences that are close to the moment the pain stimuli signal changes its intensity. We synchronize all the modalities to have the same sampling rate with the video.
Preprocessing. To compute the latent representation for each video frame in the dataset, we use EMOCA [19] to extract the expression code $\psi$ and jaw pose $\theta_{\text{jaw}}$, while calculating the mean face for all other features. To address the variance difference of $\theta_{\text{jaw}}$ being too small compared to $\psi$, which can cause jaw jiggling during generation, we scale $\theta_{\text{jaw}}$ up by 100 times to balance the loss term and ensure that both $\psi$ and $\theta$ are optimized more evenly. To determine the pain expression configuration for each subject, we calculate the Prkachin and Solomon Pain Intensity (PSPI) [38] by extracting action units (AUs) using the state-of-the-art GraphAU [39]. We then compute the average PSPI for each identity, referring to this as the pain expressiveness configuration. Additionally, we extract the emotion index for each subject using the HSEmotion emotion extractor [40], averaging across frames to create a consistent emotion configuration for each subject.
In its final form, the pain facial expression dataset consists of 4 modalities: facial latent vector, pain stimuli signal, pain expressiveness configuration, and emotion configuration. Our model is trained to generate the facial latent vector conditioned on the other modalities.
@@ IV-B Experiments setup
Metrics. To compare the effectiveness of expressing pain, we draw from metrics used in multiple facial reaction generation [18, 6] and human behavior generation [13, 41, 42, 43] to propose the following set of evaluation metrics:
  • PainSim: Uses Dynamic Time Warping (DTW) to measure the temporal signal similarity between the generated sequence’s PSPI signal and the ground truth PSPI signal under the same pain stimuli.
  • PainCorr: Uses the Pearson Correlation Coefficient (PCC) to quantify the linear correlation between the generated PSPI signal and the ground truth PSPI signal.
  • PainAcc: Uses DTW to measure the similarity between the generated PSPI signal and the pain stimuli signal, verifying the relationship between the stimuli and the expression.
  • PainDist: Uses Pairwise Mean Squared Error (MSE) to evaluate the difference between the generated expressions and the ground truth expressions.
  • PainDivrs: MSE of multiple generated expressions under the same stimuli signals to assess the diversity of the generated outputs.
  • PainVar: Measures the variance of generated expressions within the same sequence to evaluate how varied the expressions are in a single sequence.
Baselines. We establish three baselines to validate the model’s effectiveness:
@@
  • Vector Quantized VAE and Autoregressive Model: We test the effectiveness of the diffusion model compared to common approaches by applying the SOTA method in facial reaction generation from [44] on this dataset for comparison. For short, we refer to this method as autoregressive.
Implementation details. We train PainDiffusion with a sequence length of 64, a warm-up phase of 5k steps, and a total of 300k training steps. The model uses a frame stack of 4, 35 sampling steps, a learning rate of $4\times 10^{-4}$, and an exponential moving average (EMA) with a decay of 0.999. Training is conducted on a pair of NVIDIA 4060Ti and 3080 GPUs. All the metrics and qualitative output are computed with a generation length of 640 frames, 10 times longer than the training horizon, to confirm the model’s ability to generate arbitrary-length output.

### v2

# PainDiffusion: Learning to Express Pain

@@ Abstract
Accurate pain expression synthesis is essential for improving clinical training and human-robot interaction. Current Robotic Patient Simulators (RPSs) lack realistic pain facial expressions, limiting their effectiveness in medical training. In this work, we introduce PainDiffusion, a generative model that synthesizes naturalistic facial pain expressions. Unlike traditional heuristic or autoregressive methods, PainDiffusion operates in a continuous latent space, ensuring smoother and more natural facial motion while supporting indefinite-length generation via diffusion forcing. Our approach incorporates intrinsic characteristics such as pain expressiveness and emotion, allowing for personalized and controllable pain expression synthesis. We train and evaluate our model using the BioVid HeatPain Database. Additionally, we integrate PainDiffusion into a robotic system to assess its applicability in real-time rehabilitation exercises. Qualitative studies with clinicians reveal that PainDiffusion produces realistic pain expressions, with a 31.2% ± 4.8% preference rate against ground-truth recordings. Our results suggest that PainDiffusion can serve as a viable alternative to real patients in clinical training and simulation, bridging the gap between synthetic and naturalistic pain expression. Code and videos are available at: https://damtien444.github.io/paindf/.
@@ I INTRODUCTION
Reading patient pain accurately is crucial for improving clinical care [1]. However, research indicates that clinicians often underestimate or misinterpret patient pain compared to laypeople, possibly due to cognitive biases or overreliance on medical instruments [2, 3]. This underestimation can lead to inadequate pain management, misdiagnosis, and patient distress, increasing health risks [4].
Pain is a multimodal phenomenon involving biological signals, facial expressions, heart rate, skin color changes, speech tone, and more [5]. Among these, facial expressions provide critical nonverbal cues that help clinicians assess pain intensity and emotional distress. Training clinicians to recognize these expressions more effectively could improve patient care [6].
Robotic Patient Simulators (RPSs) allow healthcare professionals to practice procedures and diagnostic skills without risking patient harm. Current RPSs can simulate limb movements, breathing, bleeding, and biosignals, but they often lack realistic facial expressions [7]. Given that 70% of medical errors stem from communication issues, and 75% of those lead to patient’s death [8], enhancing RPSs with realistic facial reactions could significantly improve that communication.
@@
Figure 1: Overview. PainDiffusion inputs pain stimuli signals, expressiveness configuration, emotion status, and past frames to generate the next appropriate pain facial reaction.
A major challenge in incorporating facial pain expressions into RPSs is the high-dimensional, nonlinear mapping between pain stimuli and facial responses. Pain expressions are inherently probabilistic and modulated by inter-individual factors such as demographic attributes, baseline expressivity, emotional state, and pain type. Traditional heuristic or rule-based systems fail to generalize across such complexity, as they rely on static, predefined mappings that do not capture the stochastic nature of facial responses [6, 9, 10].
We hypothesize that leveraging a naturalistic, non-acted pain dataset can better model the variability and uncertainty inherent in pain expressions. Recent advancements in deep generative models, including diffusion models, variational autoencoders (VAEs), offer a principled approach for learning the latent structure of facial pain expressions. By conditioning generative models on multimodal pain-related signals (e.g., physiological markers, stimulus properties, and affective states), we can develop a data-driven system that synthesizes realistic facial expressions in response to dynamic pain stimuli. This approach bypasses the limitations of manually designed animation heuristics, instead enabling adaptive, personalized facial expression generation that aligns with real-world pain perception dynamics.
In this work, we focus on synthesizing facial reactions based on embodiment signals. We exclusively use the BioVid HeatPain Database [11], which is the only dataset that includes direct recordings of stimuli, physiological signals, and naturalistic facial reactions. This restricts our ability to generalize across diverse patient populations and cultural backgrounds. Our contributions are threefold: (1) We introduce PainDiffusion, a model designed to generate pain-related facial expressions with arbitrary-length predictions, making it suitable for robotic applications. It incorporates intrinsic characteristics such as expressiveness and emotion, allowing for more controllable and personalized generation. (2) We propose a new set of baselines and metrics to effectively evaluate the quality and accuracy of pain expressions generated by our model. (3) Finally, we integrate PainDiffusion with a robotic elbow and have rehabilitation clinicians assess the pain reactions to their actions.
@@ II RELATED WORKS
In training robots, earlier work in pain synthesizing primarily focused on recognizing pain situations and selecting expressions from a predefined set [10, 11]. Among pain-related research, most efforts have focused on recognition and classification using either traditional machine learning or deep learning techniques [12, 9, 6]. However, those methods often results in expressions that are unnatural and not automatic. More recently, research has shifted towards using generative models for facial expressions [13]. Meanwhile, diffusion models and score-based models have emerged as powerful tools for generating images and videos. These models have now achieved state-of-the-art performance, particularly in generating realistic images and videos [14, 15, 16]. In human action generation, recent work is increasingly focusing on diffusion models to achieve state-of-the-art results [17, 18, 19].
We ultimately chose diffusion models for four key reasons. First, they operate in a continuous data domain, enabling smoother and more natural facial motion—an area where autoregressive models struggle [19, 14]. Second, diffusion models support diffusion forcing, allowing for indefinite-length signal generation without divergence, a challenge faced by both GAN-based and autoregressive approaches [20]. Third, they have demonstrated high-quality performance across multiple modalities as mentioned previously, making them a robust choice for pain expression synthesis. Finally, diffusion models offer controllability over the influence of different conditioning signals through classifier-free guidance, enhancing their adaptability to diverse use cases [21].
@@ III PAIN DIFFUSION
Our main goal is to model the relationship between facial expressions, pain-causing signals, and the intrinsic features of an individual. To achieve this, we define the high-level task as follows: Given a continuous sequence of pain-causing signals and the configuration of the individual, we autoregressively predict the appropriate facial expression.
@@ III-C Temporal Latent Unet
Latent diffusion models (LDMs) have been effectively utilized for generating images, videos, and human behaviors [26, 19, 14, 27]. In our work, we follow the LDM approach by restricting data representation to the EMOCA latent space, resulting in a much smaller data space compared to conventional diffusion models for image or video generation. Prior works in video generation diffusion models [28, 26] have introduced temporal layers to the standard spatial U-Net [29] to better capture temporal information. Instead of employing 2D convolutions as in [16], we hypothesize that our facial latent space does not retain substantial structural information of $I$ after multiple layers of encoding in EMOCA [22]. Therefore, using 1D convolutions is sufficiently effective.
The network is designed with convolutional ResNet blocks [30], each followed by a spatial attention block and then a temporal attention block. The noise $\lambda_{t}$ is integrated into the ResNet block using scale and shift operators, while conditioning information $c=(C,\mathcal{P},\mathcal{E})$ is encoded with a lightweight MLP encoder and incorporated into the spatial attention block via cross-attention. To embed temporal information, we scale and shift the features using temporal embeddings. Both the temporal and noise-time embeddings are encoded using sinusoidal embeddings for positional information. The up and down blocks have the architecture as illustrated in Fig. 2.
@@ III-D Elucidated Diffusion
where $\boldsymbol{z}$ is the facial latent vector, $\sigma$ is the standard deviation of the Gaussian noise level, $C$ represents conditions, and $F_{\theta}$ is the temporal latent U-Net. By adjusting the terms $c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$, different diffusion strategies can be achieved with minimal changes to the model. We adopt the $c$ terms from the Elucidated Diffusion Model (EDM) [31] due to the flexibility of the network architecture. As a result, we do not consider reparameterization approaches, since EDM [31] is a hybrid of both velocity, start, and noise prediction.
For clarity, with $y$ being the clear sample and $n$ being noise, the training objective is simplified as:
@@
Figure 3: A clinician performing an elbow range-of-motion rehabilitation exercise while observing the virtual avatar’s reaction.
@@ III-E Diffusion Forcing
The goal of this model is to be applied to a robot and generate arbitrary-length predictions. We adopt diffusion forcing [20] to extend beyond the short training horizon. Diffusion forcing assigns different noise levels to each temporal timestep, placing more uncertainty on future frames while reducing uncertainty for past frames. It also introduces a hyperparameter that controls uncertainty—higher values result in greater uncertainty, which in turn requires more denoising steps.
@@
III-F Elbow Range of Movement Exercise Robot
We build a simple single-joint elbow robot using an MX-64 motor controlled by an Arduino. The joint’s angular range is linearly mapped to the heat stimulus values from the BioVid HeatPain Database, simulating a patient experiencing pain when flexing their elbow. To create a realistic facial representation, we employ Gaussian Avatars [33] to map the 3D FLAME mesh [23] onto a lifelike avatar. This setup is used to allow clinicians to assess the realism of the real-time rehabilitation exercise.
@@ IV EXPERIMENTS
Our experiments are designed with two primary goals. First, we aim to demonstrate that our model outperforms the baseline and common approaches in facial expression generation. Second, we seek to show that our model can produce arbitrary-length predictions without divergence. To achieve these objectives, we propose a new set of metrics specifically for evaluating pain facial expressions and long-term prediction errors. We then conduct a human evaluation with both laypersons and clinicians to assess the model’s naturalistic.
@@ IV-A Dataset
To create the pain facial expression dataset, we use the BioVid Heatpain Database part C [11], which captures heat pain responses from 87 subjects with 4 levels of pain intensity and 30 minutes of data, separated by pauses. The original dataset includes 3 modalities: frontal face video (25 FPS), biomedical signals, and heat stimuli signals. We split the dataset into two subsets: 61 subjects for training and 26 subjects for validation. The validation subsets maintain an equal ratio of male and female participants, and include 5 low-expression and 21 normal-expression subjects. We filter the original validation videos to contain sequences that are close to the moment the pain stimuli signal changes its intensity. We synchronize all the modalities to have the same sampling rate with the video.
Preprocessing. To compute the latent representation for each video frame in the dataset, we use EMOCA [22] to extract the expression code $\psi$ and jaw pose $\theta_{\text{jaw}}$, while calculating the mean face for all other features. To determine the pain expression configuration for each subject, we calculate the Prkachin and Solomon Pain Intensity (PSPI) [34] by extracting action units (AUs) using the state-of-the-art GraphAU [35]. We then compute the average PSPI for each identity, referring to this as the pain expressiveness configuration. Additionally, we extract the emotion index for each subject using the HSEmotion emotion extractor [36], averaging across frames to create a consistent emotion configuration for each subject.
In its final form, the pain facial expression dataset consists of 4 modalities: facial expression parameters, pain stimuli signal, pain expressiveness configuration, and emotion configuration. Our model is trained to generate the facial expression parameters conditioned on the other modalities.
@@ IV-B Experiments setup
Quantitative metrics. To compare the effectiveness of expressing pain, we draw from metrics used in multiple facial reaction generation [37] and human behavior generation [19] to propose the following set of evaluation metrics on two modalities of PSPI and FLAME expression parameters:
  • Sim: Dynamic Time Warping (DTW) to measure the temporal signal similarity between the generated sequence’s PSPI signal and the ground truth PSPI signal under the same pain stimuli.
  • Corr: Uses the Pearson Correlation Coefficient (PCC) to quantify the linear correlation between the generated PSPI signal and the ground truth PSPI signal.
  • Dist: Uses Pairwise Mean Squared Error (MSE) to evaluate the difference between the generated expressions and the ground truth expressions.
  • Divrs: MSE of multiple generated expressions under the same stimuli signals to assess the diversity of the generated outputs.
  • Var: Measures the variance of generated expressions within the same sequence to evaluate how varied the expressions are in a single sequence.
Baselines. We establish three baselines to validate the model’s effectiveness. Because generating :
@@
  • Vector Quantized VAE and Autoregressive Model: We use the winning model from multiple appropriate facial reactions challenge - REACT Challenge 2024111The Second REACT Challenge@IEEE FG24 [38] with modifications to take stimuli signal as input. For short, we refer to this method as autoregressive.
Figure 4: Clinicians’ ratings of the virtual avatar’s quality after performing the elbow range-of-motion exercise. A rating of 1 indicates the lowest quality, while 5 indicates the highest.
Implementation details. We train PainDiffusion with a sequence length of 64, a warm-up phase of 5k steps, and a total of 300k training steps, a learning rate of $4\times 10^{-4}$, and an exponential moving average (EMA) with a decay of 0.999. Training is conducted on a pair of NVIDIA 3080 GPUs. All the metrics and qualitative output are computed with a generation length of 640 frames, 10 times longer than the training horizon, to confirm the model’s ability to generate arbitrary-length output.
Qualitative experiment setup. To evaluate the realism of the generated pain expressions, we conduct a user study involving healthcare professionals. Specifically, we recruited 18 Japanese rehabilitation clinicians, all of whom regularly interact with patients experiencing pain. The experiment consists of two main phases: real-time interaction evaluation and video-based preference testing.
In the first phase, participants interacted with the robotic elbow for two minutes, observing the corresponding immediate facial reactions displayed on a screen. Figure 3 illustrates the experimental setup for this phase. We ran a pilot experiment with 10 laypeople to decide which questions to ask. The clinicians were then asked to complete a questionnaire assessing the realism of the generated expressions across three key dimensions: (1) Response Dynamics: ”How swift and strong is the reaction to the stimuli?”; (2) Motion Realism: ”Do the generated movements appear realistic?”; (3) Patient Resemblance: ”Does the reaction resemble that of a real patient?”. Participant ratings for each question were recorded on a 5-point Likert scale.
The second phase involves a preference test using video-based comparisons. Participants were presented with 22 questions, assessing their preferences based on three aspects: (1) Temporal consistency with stimuli signals (2 questions), (2) Realism of facial reactions against groundtruth (8 questions), (3) Diversity of facial expressions (9 questions).
The experiment is implemented using the jsPsych framework222jsPsych is a JavaScript framework for creating behavioral experiments that run in a web browser.. From a dataset of 50 random validation video samples, each question is randomly selected from the dataset. To assess diversity, we generate four variations of the same stimuli. For both evaluating temporal consistency and diversity, participants rate the strength of the reaction on a 5-point scale. Each question is separated by a fixed screen in one second to inform the user of the boundary between the questions. We learn that cropping mouth and eye regions helps users better assess diversity as it limits the cognitive load of the comparison. We have 2/3 of diversity questions in the format of cropping eye and mouth.
Figure 5: The distribution of clinicians’ rating for temporal consistency with stimuli signal of PainDiffusion and Groundtruth in the video-preference experiment.
Figure 6: Visualization of four sample stimuli from the BioVid HeatPain Database validation set, ranging from level 1 (a) to level 4 (d) (darker orange indicates higher pain intensity). We ran five independent predictions for both PainDiffusion and the autoregressive baseline and plotted their range distributions as green and gray shaded areas, respectively. The orange-shaded region represents the duration during which the pain stimulus was applied to the subjects. PainDiffusion was configured with an emotion of contempt and an expression intensity of 8.5 on the PSPI scale [34]. The ground truth PSPI signal predominantly overlaps with the green-shaded region of PainDiffusion, whereas the autoregressive baseline tends to exhibit higher pain intensity on the PSPI scale.